A certain experiment conducted was normally distributed. The mean value is 50 and the standard deviation is 4. If the total population is 1000 (discrete variable), calculate the percentage of the population within z=-1.4 and z=2.56


1
Expert's answer
2021-11-22T19:33:19-0500

"P(-1.4<Z<2.56)=P(Z<2.56)-P(Z<-1.4)="

"=0.9948-0.0808=0.9140=91.4\%."